A person jumps onto a cement floor from a height of 1 m and comes to rest in 0.1 second .the same person on jumping from a height of 9 m into a sand pit, comes to rest in 1 second . Compare the forces exerted on him by cement floor and sand pit.

Grade:9

1 Answers

Uttam Gulshan
13 years ago

h=1 m.  h'=9 m.

t=0.1 s. t'=1 s.

Now force exerted in first case be F and in second case be F'.

Let the mass of man be m.

F=(m((2gh)^(1/2)))/t

F'=(m((2gh')^(1/2)))/t'

=> F/F'=((h/h')^(1/2))*(t'/t)

=>F/F'=10/3.

Hence , F'=(3/10)F.